# Heterogeneous Multi-Agent Reinforcement Learning for Zero-Shot Scalable Collaboration

### v1

@@ VI Conclusion
In this paper, we propose a novel MARL framework SHPPO to add both inter-individual and temporal heterogeneity to any parameter-shared actor-critic architecture so that the framework can be scalable for unseen zero-shot new tasks. We introduce latent learning to adaptively update the parameters of the heterogeneous layer and the agents’ strategies. Our approach shows superior performance over baselines, especially for zero-shot scalable collaboration, on several classic MARL tasks. Future work could implement the heterogeneous layer with more complex network designs like transformers, or integrate population-invariant methods to further enhance scalability. One can even take advantage of the emerging large language models (LLMs), converting the latent variables to heterogeneous prompts for LLM agents to gain better scalability.
Appendix A Environment Settings
We modify and keep the number of observations the same in one task series, so that we can focus on the transfer of strategies without mapping the observations. The detailed settings are shown in Table IV.
TABLE IV: Number of observations in the modified tasks for scalability test.
Task #Obs of allies* #Obs of enemies
MMM series 9 10
8m_vs_9m series 5 6
3_vs_1_with_keeper series 11 11
counterattack_easy series 11 11
    The number of observations of the ally agents includes the agent itself.
Appendix B Network Settings
We provide the network configurations as follows:
TABLE V: Network configurations for SHPPO.
@@ Appendix B Network Settings
Appendix C Parameter Settings
The hyperparameters used for training of SHPPO is provided in Table VI.
TABLE VI: Hyperparameters settings for SHPPO.
@@ Appendix C Parameter Settings
entropy loss weight $\lambda_{e}$ 0.001
distance loss weight $\lambda_{d}$ 0.01
inference loss weight $\lambda_{v}$ 0.1
@@
Appendix D Biography Section
[Uncaptioned image] Xudong Guo received the B.Eng. degree in automation from Tsinghua University, Beijing, China, in 2020. He is now pursuing the Ph.D. degree at the Department of Automation, Tsinghua University, Beijing, China. His research interests include multi-agent systems and reinforcement learning.
[Uncaptioned image] Daming Shi received the B.S. degree in automatic science from Beihang University, Beijing, China, in 2018 and the Ph.D. degree in automatic science from Tsinghua University, Beijing, China, in 2023. His research interests include reinforcement learning and multi-agent systems in gaming and production.
[Uncaptioned image] Junjie Yu received the B.S. degree in automatic science from Beihang University, Beijing, China, in 2023. He is now pursuing the M.S. degree at the Department of Automation, Tsinghua University, Beijing, China. His current research interests include multi-agent systems and reinforcement learning.
[Uncaptioned image] Wenhui Fan received the Ph.D. degree in mechanical engineering from Zhejiang University, Hangzhou, China, in 1998. He obtained the postdoctoral certificate from Tsinghua University, Beijing, China, in 2000. He is a vice president of China Simulation Federation. He is currently a professor at Tsinghua University, Beijing, China. His current research interests include multi-agent modeling and simulation, large scale agent modeling and simulation, and multi-agent reinforcement learning.

### v2

@@ VI Conclusion
In this paper, we propose a novel MARL framework Scalable and Heterogeneous Proximal Policy Optimization (SHPPO) to integrate both inter-individual and temporal heterogeneity to any parameter-shared MARL backbone, so that the framework can learn adaptive heterogeneous strategies to achieve zero-shot scalable collaboration. We introduce latent learning to adaptively update the parameters of the heterogeneous layer and the agents’ strategies according to the transferred scales. Our approach shows superior performance on both original tasks and scalability tests over baselines in several classic MARL tasks. Future work could implement the heterogeneous layer with more complex network designs like transformers, or integrate population-invariant methods to further enhance scalability. One can even take advantage of the emerging large language models (LLMs), converting the latent variables to heterogeneous prompts for LLM agents to gain better scalability.
Appendix A Technical Appendix
Environment settings. We modify and keep the number of observations the same in one task series, so that we can focus on the transfer of strategies without mapping the observations. For example, in the task MMM2 of SAMC, there are 10 controlled agents and 12 enemies, respectively. In the original tasks of SMAC, the agent can observe all the enemies, resulting in easy and unscalable tasks. We modify the tasks to limit the observation range to the closest 10 enemies and the closest 8 allies so that we can create a series of new tasks with different numbers of agents and enemies to test scalability. Similarly, in 8m_vs_9m, we limit the range to the closest 6 enemies and the closest 5 allies. While in GRF, there are always 11 players in one team. We change the number of controlled players to test scalability but keep the observations as the global observations of all the players in the field.
Training details. We provide the network configurations in Table IV and hyperparameters used for training in Table V. For each task, we train SHPPO on one single NVIDIA V100 GPU. We conduct 20 million training steps for the SMAC environments and 25 million training steps for the GRF environments. The training time is shown in Table VI.
TABLE IV: Network configurations for SHPPO.
@@ Appendix A Technical Appendix
TABLE V: Hyperparameters settings for SHPPO.
@@
entropy loss weight $\lambda_{e}$ 0.01
distance loss weight $\lambda_{d}$ 0.1
@@
TABLE VI: Training Time of SHPPO.
Task Time (h)
MMM2 25.5$\pm$0.2
8m_vs_9m 20.3$\pm$0.6
3_vs_1_with_keeper 26.1$\pm$0.8
counterattack_easy 15.4$\pm$0.5